OpenStudy (anonymous):

find each product. (x+5)(x-5) explain

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

use the foil method:

OpenStudy (alexandervonhumboldt2):

x^2-25

OpenStudy (alexandervonhumboldt2):

(a-b)*(a+b)=a^2-b^2

OpenStudy (anonymous):

yeah its foil lol

OpenStudy (anonymous):

please don't just provide answers, let the asker PARTICIPATE in solving their own problems . If you haven't read the Code of Conduct here's the link: http://openstudy.com/code-of-conduct . I'm telling you this because I have been warned plenty times because of providing answers (:

OpenStudy (anonymous):

im bad at math lol, and i need to pass math this year

OpenStudy (anonymous):

so how would i solve it like work it out?

OpenStudy (anonymous):

because i need to show work

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{example\\ (x+1)(x-1)\\ =x\times x-x+x-1\\ =x^2-1 }\end{align}\)

OpenStudy (mathmath333):